# Selection of Filters for Photonic Crystal Spectrometer Using Domain-Aware Evolutionary Algorithms

### v1

@@ 18       $\mathbf{x}^{*}\leftarrow\mathbf{y}^{(1)}$;
Algorithm 2 Proposed DD-$(\mu/2+\lambda)$ EA, equipped with a distance-driven heuristic for integer programming, where objective function $F:\mathbb{L}^{M}\to\mathbb{R}$ is minimized. Points in its domain have dimensionality $M$. Value at every component of every point is an integer from the set $\mathbb{L}\coloneqq\overline{1,L}$, which is a metric space with distance $d_{\mathbb{L}}$. The maximal number of $F$ evaluations is limited by constant $b$. On top of that, the algorithm uses integer constants $\mu,\lambda,R$ and a real constant $0\leq m\leq 1$.
@@
Then algorithm runs the main optimization loop in line 2. EA iteratively makes changes to the candidate solutions and selects more promising candidates until the budget $b$ is exhausted. The value $b$ limits the number of objective function evaluations that the algorithm is allowed to make to find an estimate of the optimized function minimizer. We used $\widehat{\mathrm{F}}_{K}$ as the objective function with fixed $K=10^{3}.$ It is possible to use the values of the underlying random variables to distinguish statistically significant differences in the objective function from insignificant ones. However, we leave this for future work.
@@
Changes to the selected solutions are made in lines 2, 2. In line 2 algorithm performs uniform crossover. In line 2, the distance-driven mutation is applied. The details of the function Mutation are discussed in Section 5.2.2. The “strength” of the mutation is defined by the number $s\in(0,1)$ sampled from $\mathcal{D}$ in line 2. This number is transformed to the distance in $\mathbb{M}(d_{\mathbb{L}})$ in line 2. The formula for this transformation is explained in Section 5.2.1.
@@
$\delta_{\text{min}}(d_{\mathbb{L}},\bm{x})\coloneqq\min\limits_{\bm{y}\in\mathbb{L}^{M}}\left\{\textsc{LAP}(d_{\mathbb{L}},\bm{x},\bm{y})\,\middle|\,\bm{x}\not\equiv\bm{y}\right\}\\ =\min\limits_{\begin{subarray}{c}y\in\mathbb{L},\\ i\in\overline{1,M}\end{subarray}}\left\{d_{\mathbb{L}}(x_{i},y)\,\middle|\,x_{i}\neq y\right\}.$
@@
$\delta^{**}(d_{\mathbb{L}})\coloneqq\max\limits_{\bm{x}\in\mathbb{L}^{M}}\left\{\delta_{\text{min}}(d_{\mathbb{L}},\bm{x})\right\}\\ =\max\limits_{x\in\mathbb{L}}\left\{\min\limits_{y\in\mathbb{L}}\left\{d_{\mathbb{L}}(x,y)\,\middle|\,x\neq y\right\}\right\}.$
@@
The biggest distance $\delta_{\text{max}}(d_{\mathbb{L}})$ on $\mathbb{L}^{M}$ is defined analogically to the smallest: $\delta_{\text{max}}(d_{\mathbb{L}})=\max\left\{\textsc{LAP}(d_{\mathbb{L}},\bm{x},\bm{y})\,\middle|\,\bm{x},\bm{y}\in\mathbb{L}^{M}\right\}.$ We did not observe the same properties for the maximal distance as for the minimal. Therefore, we applied the general procedure proposed in [61] to compute $\widetilde{\delta_{\text{max}}}\!\left(d_{\mathbb{L}},\mathbf{x}^{(i)}\right),i\in\overline{1,R}$ for $R=10$ points chosen uniformly at random from $\mathbb{L}^{M}.$ The maximal of this values is used as $\widehat{\delta_{\text{max}}}(d_{\mathbb{L}})$:
@@
[…]
@@
[…]
1 Initialization of parent: $\mathbf{x}^{(1)},\mathrm{v}^{(1)}\leftarrow\bm{x}^{(0)},0$;
@@ 27                   $\pi^{(i+1)}\leftarrow\pi_{\mathbf{y}}$;
Algorithm 3 Proposed method to locate a vector $\bm{y}\in\mathbb{L}^{M}$ such that $\textsc{LAP}(d_{\mathbb{L}},\bm{y},\bm{x}^{(0)})\approx s$, given $M\in\mathbb{N}$, $\bm{x}^{(0)}\in\mathbb{L}^{M}$, constant $s$ and metric $d_{\mathbb{L}}$ on the set $\mathbb{L}\coloneqq\overline{1,L}$. The maximal number of LAP function evaluations is limited by the constant $b$. On top of that, the algorithm uses constants $\lambda,R\in\mathbb{N}$.
[…]
The general scheme of Algorithm 3 is the same as in $(1+\lambda)$ EA. It treats $|\textsc{LAP}(d_{\mathbb{L}},\bm{x},\bm{y})-s|$ as the minimized objective function in the search space $\mathbb{L}^{M}$ and iteratively produces $\lambda$ candidate solutions by making variations of best-so-far solution. However, this algorithm takes advantage of the permutation $\pi$ available through $\textsc{LAP}^{*}$. The maintained bijection $\pi^{(i)}$ represents the assignment of components in the initial solution $\bm{x}^{(0)}$ and the current best-so-far $\mathbf{x}^{(i)}.$ In the variation phase Algorithm 3 in line 3 selects a component with number $\mathrm{k}$ in $\bm{x}^{(0)}$. The filter $x^{(0)}_{\mathrm{k}}$ at position $\mathrm{k}$ corresponds to some component in the $\mathbf{x}^{(i)}$ through the permutation $\pi^{(i)}$ as shown in Figure 8.
@@
The key idea of our heuristic is to increase the distance $d_{\mathbb{L}}(x^{(0)}_{\mathrm{k}},\mathrm{x}^{(i)}_{\pi^{(i)}(\mathrm{k})})$ between those filters, when $\textsc{LAP}(d_{\mathbb{L}},\bm{x}^{(0)},\mathbf{x}^{(i)})<s$ and reduce otherwise. This is implemented in lines 3, 3, 3, 3, 3 and 3, that generate candidate solution $\mathbf{y}$ with the described change in one of its components. The implementation details from the pseudocode are visualized in Figure 9. For the sake of shortness in our pseudocode Algorithm 2 and Figure 9 we represent $\textsc{LAP}(d_{\mathbb{L}},\bm{x}^{(0)},\mathbf{x}^{(i)})$ by $\mathrm{v}^{(i)}$, and $\textsc{LAP}(d_{\mathbb{L}},\bm{x}^{(0)},\mathbf{y})$ by $\mathrm{v}_{\mathbf{y}}$. It is not guaranteed that $\mathrm{v}_{\mathbf{y}}$ is smaller or bigger than $\mathrm{v}^{(i)}$ because the assignment $\pi_{\mathbf{y}}$ might be not equal to $\pi^{(i)}$. Therefore, we randomize this heuristic and use it as a variation operator in the inherent structure of EA.
@@
The choice of the new value for the considered component is done through Discrete Harmonic Distribution in lines 3 and 3 of Algorithm 3. It is defined and applied for mutations for example in [39], where it is shown that sampling from it gives the best possible number of iterations to locate the optimum of the problem considered in that work. Our choice of this distribution is motivated by that rigorous result. The probabilities assigned to integers via such harmonic sampling are visualized in lower part of Figure 9. Notably, the described variation may not change the candidate solution, therefore, we repeat the procedure $R$ times unless $\mathbf{y}$ becomes different from $\mathbf{x}^{(i)}$ earlier.
@@
Numbers in this family of sets represent the relative deviation from the target distance and the distance between the mutant and the original point. The smaller those values are, the smaller the error is, and so the more precise our mutation operator is. We observed that for each $i\in\overline{1,100}$ the mean $m^{(i)}$ of $\mathbb{V}^{(i)}$ was smaller than $0.05$ for both $d_{1}$ and $d_{2}$ taken as $d_{\mathbb{L}}$. The median of $\left\{m^{(i)}\right\}_{i=1}^{100}$ is $\approx 4\cdot 10^{-5}$ for $d_{1}$ and $\approx 10^{-4}$ for $d_{2}$.
@@ 5.3 Comparison of the Defined Metrics
(b) $\textsc{LAP}\!\left(d_{1},\mathbf{x}^{(i)},\mathbf{y}^{(i,j)}\right)\in(10\cdot\delta_{1},15\cdot\delta_{1}),$
rejected $44\%$
(c) $\textsc{LAP}\!\left(d_{2},\mathbf{x}^{(i)},\mathbf{y}^{(i,j)}\right)\in(10\cdot\delta_{2},15\cdot\delta_{2}),$
rejected $15\%$
@@
The neighborhoods are defined as a set of candidate solutions located close to each other. It is likely to obtain one of the neighbors when making a small mutation. We will use Hamming distance to describe the conventional mutation in $(\mu+\lambda)$ EA and $(\mu/2+\lambda)$ EA. This metric is defined for any $\bm{x},\bm{y}\in\mathbb{L}^{M}$ as:
@@
Since it is not known which distance is better for OFS, it is relevant to make their comparison. In this section, we propose the first step towards a comprehensive understanding of their difference. Specifically, we analyze how many solutions with statistically different values of $\widehat{\mathrm{F}}_{K}$ are located in neighborhoods for the three cases of inherent metric: 1) Hamming distance $\mathscr{H}$, 2) $\textsc{LAP}(d_{1},\cdot,\cdot)$, 3) $\textsc{LAP}(d_{2},\cdot,\cdot)$. We designed the experiment outlined in Algorithm 4 to perform this analysis.
@@ 16             $\mathrm{P}_{i,j}\leftarrow\textsc{Welch}\!\left(\mathbf{w}^{(i)},\mathbf{z}^{(i,j)}\right)$;
Eventually, Algorithm 4 returns all parents, corresponding mutants, and the matrix $\mathbf{P}$. This matrix is visualized in Figure 11 for the three cases of inherent distance: $\mathscr{H},\textsc{LAP}(d_{1},\cdot,\cdot),\textsc{LAP}(d_{2},\cdot,\cdot)$. When one statistical comparison is made, it is natural to reject the null hypothesis when pvalue is smaller than $0.05$. Since the goal of the experiment is to draw conclusions on the number of statistically different neighbors, we take into account Bonferroni correction. That is, we define the maximal pvalue that is still acceptable to reject the null hypothesis by $n^{2}=1024$. So, when the pvalue is smaller than $0.05/1024$, the corresponding points are considered different and the square is colored in blue in Figure 11. Otherwise, the null hypothesis is not rejected and the square is colored in red.
@@ 5.4 Distance-Driven UMDA
1 For every $j\in\overline{1,L}$ initialize: $\mathrm{p}^{(1)}_{j}\leftarrow 1/L$;

### v2

@@ 18 $\mathbf{x}^{*}\leftarrow\mathbf{y}^{(1)}$;
Algorithm 2 Proposed $D$-$(\mu/2+\lambda)$ EA, equipped with a distance-driven heuristic for integer programming, where objective function $F:\mathbb{L}^{M}\to\mathbb{R}$ is minimized. Points in its domain have dimensionality $M$. Value at every component of every point is an integer from the set $\mathbb{L}\coloneqq\overline{1,L}$, which is a metric space with distance $d_{\mathbb{L}}$. The maximal number of $F$ evaluations is limited by constant $b$. On top of that, the algorithm uses integer constants $\mu,\lambda,R$ and a real constant $0\leq m\leq 1$.
@@
The algorithm then runs the main optimization loop in line 2. EA iteratively makes changes to the candidate solutions and selects more promising candidates until the budget $b$ is exhausted. The value $b$ limits the number of objective function evaluations that the algorithm is allowed to make to find an estimate of the optimized function minimizer. We used $\widehat{\mathrm{F}}_{K}$ as the objective function with fixed $K=10^{3}.$
We also experimented with application of Welch’s T-test to the values of the underlying random variables to distinguish statistically significant differences in the objective function from insignificant ones. This application did not lead to significant improvement, and in some cases it even worsened performance of the corresponding algorithms. Therefore, we leave application of statistical selection methods for future work.
@@
Changes to the selected solutions are made in lines 22. In line 2 algorithm performs uniform crossover. In line 2, the distance-driven mutation is applied. The details of the function Mutation are discussed in Section 5.3.2. The “strength” of the mutation is defined by the number $s\in(0,1)$ sampled from $\mathcal{D}$ in line 2. This number is transformed to the distance in $\mathbb{M}(d_{\mathbb{L}})$ in line 2. The formula for this transformation is explained in Section 5.3.1.
@@
$\delta_{\text{min}}(d_{\mathbb{L}},\bm{x})\coloneqq\min\limits_{\bm{y}\in\mathbb{L}^{M}}\left\{\textsc{LAP}(d_{\mathbb{L}},\bm{x},\bm{y})\,\middle|\,\bm{x}\not\equiv\bm{y}\right\}=\min\limits_{\begin{subarray}{c}y\in\mathbb{L},\\ i\in\overline{1,M}\end{subarray}}\left\{d_{\mathbb{L}}(x_{i},y)\,\middle|\,x_{i}\neq y\right\}.$
@@
$\delta^{**}(d_{\mathbb{L}})\coloneqq\max\limits_{\bm{x}\in\mathbb{L}^{M}}\left\{\delta_{\text{min}}(d_{\mathbb{L}},\bm{x})\right\}=\max\limits_{x\in\mathbb{L}}\left\{\min\limits_{y\in\mathbb{L}}\left\{d_{\mathbb{L}}(x,y)\,\middle|\,x\neq y\right\}\right\}.$
@@
The biggest distance $\delta_{\text{max}}(d_{\mathbb{L}})$ on $\mathbb{L}^{M}$ is defined by analogy to the smallest: $\delta_{\text{max}}(d_{\mathbb{L}})=\max\left\{\textsc{LAP}(d_{\mathbb{L}},\bm{x},\bm{y})\,\middle|\,\bm{x},\bm{y}\in\mathbb{L}^{M}\right\}.$ We did not observe the same properties for the maximal distance as for the minimal. Therefore, we applied the general procedure proposed in [57] to compute $\widetilde{\delta_{\text{max}}}\!\left(d_{\mathbb{L}},\mathbf{x}^{(i)}\right),i\in\overline{1,R}$ for $R=10$ points chosen uniformly at random from $\mathbb{L}^{M}.$ The maximal of this values is used as $\widehat{\delta_{\text{max}}}(d_{\mathbb{L}})$:
@@
In line 2 of Algorithm 2 we denoted the method that takes $R$ points from $\mathbb{L}^{M}$ and computes values $\widehat{\delta_{\text{min}}}(d_{\mathbb{L}}),\widehat{\delta_{\text{max}}}(d_{\mathbb{L}}),\gamma$ as $\textsc{Explore}.$ It returns only $\gamma$ since it is used later in the algorithm for the transformation of step size to distance in line 2.
@@
We propose a heuristic approach that aims at locating $\bm{y}$ to minimize $|\textsc{LAP}(d_{\mathbb{L}},\bm{x},\bm{y})-S|,$ within the limited number of LAP function evaluations. In line 2 of Algorithm 2 this method is called $\textsc{DDMutation},$ which stands for distance-driven mutation. This non-trivial internal optimization task can be seen as an inverse linear assignment problem because the goal is to produce a vector that gives a specific solution to this problem. Algorithm 3 summarizes the proposed method to solve this problem, called Distance‑Driven Assignment Evolutionary Algorithm (DDA‑EA).
The number of LAP function evaluations is restricted since it might be a computationally expensive procedure, especially for bigger values of $M$. In contrast to LAP, which returns only one real value, we use notation $\textsc{LAP}^{*}(d_{\mathbb{L}},\bm{x},\bm{y})$ to emphasize that it returns two objects: 1) the permutation $\pi^{*}$ on which the minimum of $\sum_{i=1}^{M}{d_{\mathbb{L}}(x_{i},y_{\pi^{*}(i)})}$ is reached, and 2) the inherent smallest value, achievable through the permutation $\pi^{*}.$ Notably, the second number of $\textsc{LAP}^{*}$ equals to the one returned by $\textsc{LAP}.$
The general scheme of Algorithm 3 is the same as in $(1+\lambda)$ EA. It treats $|\textsc{LAP}(d_{\mathbb{L}},\bm{x},\bm{y})-s|$ as the minimized objective function in the search space $\mathbb{L}^{M}$ and iteratively produces $\lambda$ candidate solutions by making variations of the best-so-far solution. However, this algorithm takes advantage of the permutation $\pi$ available through $\textsc{LAP}^{*}$. The maintained bijection $\pi^{(i)}$ represents the assignment of components in the initial solution $\bm{x}^{(0)}$ and the current best-so-far $\mathbf{x}^{(i)}.$ In the variation phase Algorithm 3 in line 3 selects a component with number $\mathrm{k}$ in $\bm{x}^{(0)}$.
1Initialization of parent: $\mathbf{x}^{(1)},\mathrm{v}^{(1)}\leftarrow\bm{x}^{(0)},0$;
@@ 27       $\pi^{(i+1)}\leftarrow\pi_{\mathbf{y}}$;
Algorithm 3 Distance‑Driven Assignment Evolutionary Algorithm (DDA‑EA) – the proposed method to locate a vector $\bm{y}\in\mathbb{L}^{M}$ such that $\textsc{LAP}(d_{\mathbb{L}},\bm{y},\bm{x}^{(0)})\approx s$, given $M\in\mathbb{N}$, $\bm{x}^{(0)}\in\mathbb{L}^{M}$, constant $s$ and metric $d_{\mathbb{L}}$ on the set $\mathbb{L}\coloneqq\overline{1,L}$. The maximal number of LAP function evaluations is limited by the constant $b$. On top of that, the algorithm uses constants $\lambda,R\in\mathbb{N}$. The method is used to solve the internal optimization problem for distance-driven algorithm applied to OFS.
Figure 5 shows the initial (reference) solution, $\bm{x}^{(0)}$, and the mutant solutions, $\mathbf{x}^{(i)}$, at a curtain stage of Algorithm 3. At each stage, the algorithm samples a position $k$, and finds a position in the mutant solution that corresponds to position $k$ in the initial solution, under the current permutation $\pi^{(i)}$. Thus, the filter $x^{(0)}_{k}$ at position $k$ corresponds to a component in $\mathbf{x}^{(i)}$ through the permutation $\pi^{(i)}$. The fact that the position of the mutant solution may not be the same as the position in the reference solution is indicated by the curved red line.
@@
The key idea of our heuristic is to increase the distance $d_{\mathbb{L}}(x^{(0)}_{\mathrm{k}},\mathrm{x}^{(i)}_{\pi^{(i)}(\mathrm{k})})$ between those filters, when $\textsc{LAP}(d_{\mathbb{L}},\bm{x}^{(0)},\mathbf{x}^{(i)})<s$ and reduce otherwise. This is implemented in lines 33333 and 3, that generate candidate solution $\mathbf{y}$ with the described change in one of its components. The implementation details from the pseudocode are visualized in Figure 6. For the sake of shortness in our pseudocode Algorithm 2 and Figure 6 we represent $\textsc{LAP}(d_{\mathbb{L}},\bm{x}^{(0)},\mathbf{x}^{(i)})$ by $\mathrm{v}^{(i)}$, and $\textsc{LAP}(d_{\mathbb{L}},\bm{x}^{(0)},\mathbf{y})$ by $\mathrm{v}_{\mathbf{y}}$.
In Figure 6, we demonstrate the selection procedure of the new filter in both cases of how the total distance LAP compares with $s$. When this distance is bigger than $s$, the filters to the left are considered. Otherwise, filters to the right of $f_{p}$ are considered. The selection of a filter from the considered ones is done by chance according to the Harmonic distribution [35]. This distribution selects filters closer to the initial one with higher probability than selecting those that are further. It is not guaranteed that $\mathrm{v}_{\mathbf{y}}$ is smaller or bigger than $\mathrm{v}^{(i)}$ because the assignment $\pi_{\mathbf{y}}$ might be not equal to $\pi^{(i)}$. Therefore, we randomize this heuristic and use it as a variation operator in the inherent structure of EA.
@@
The choice of the new value for the considered component is done through Discrete Harmonic Distribution in lines 3 and 3 of Algorithm 3. It is defined and applied for mutations, for example, in [35], where it is shown that sampling from it gives the best possible number of iterations to locate the optimum of the problem considered in that work. Our choice of this distribution is motivated by this rigorous result. The probabilities assigned to integers via such harmonic sampling are visualized in lower part of Figure 6. Notably, the described variation may not change the candidate solution, therefore, we repeat the procedure $R$ times unless $\mathbf{y}$ becomes different from $\mathbf{x}^{(i)}$ earlier.
@@
Numbers in this family of sets represent the relative deviation from the target distance and the distance between the mutant and the original point. The smaller those values are, the smaller the gas retrieval error is, and so the more precise our mutation operator is. We observed that for each $i\in\overline{1,100}$ the mean $m^{(i)}$ of $\mathbb{V}^{(i)}$ was smaller than $0.05$ for both $d_{1}$ and $d_{2}$ taken as $d_{\mathbb{L}}$. The median of $\left\{m^{(i)}\right\}_{i=1}^{100}$ is $\approx 4\cdot 10^{-5}$ for $d_{1}$ and $\approx 10^{-4}$ for $d_{2}$.
@@ 5.4 Comparison of the Defined Metrics
(b) $\textsc{LAP}\!\left(d_{1},\mathbf{x}^{(i)},\mathbf{y}^{(i,j)}\right)\in(10\cdot\delta_{1},15\cdot\delta_{1}),$ rejected $44\%$ (c) $\textsc{LAP}\!\left(d_{2},\mathbf{x}^{(i)},\mathbf{y}^{(i,j)}\right)\in(10\cdot\delta_{2},15\cdot\delta_{2}),$ rejected $15\%$
@@
Having defined several distance metrics to quantify differences between solutions, we now turn to comparing the local structure of the search space induced by each metric. We begin with the Hamming distance, which underlies the conventional mutation operator in the $(\mu+\lambda)$ EA and the $(\mu/2+\lambda)$ EA. This metric is defined for any $\bm{x},\bm{y}\in\mathbb{L}^{M}$ as:
@@
Since it is not known which distance is better for OFS, it is relevant to make their comparison. In this section, we propose the first step towards a comprehensive understanding of difference between metrics. Specifically, we analyze how many solutions with statistically different values of $\widehat{\mathrm{F}}_{K}$ are located in neighborhoods for the three cases of inherent metric: 1) Hamming distance $\mathscr{H}$, 2) $\textsc{LAP}(d_{1},\cdot,\cdot)$, 3) $\textsc{LAP}(d_{2},\cdot,\cdot)$. We designed the experiment outlined in Algorithm 4 to perform this analysis.
@@ 16    $\mathrm{P}_{i,j}\leftarrow\textsc{Welch}\!\left(\mathbf{w}^{(i)},\mathbf{z}^{(i,j)}\right)$;
Eventually, Algorithm 4 returns all parents, corresponding mutants, and the matrix $\mathbf{P}$. This matrix is visualized in Figure 8 for the three cases of inherent distance: $\mathscr{H},\textsc{LAP}(d_{1},\cdot,\cdot),\textsc{LAP}(d_{2},\cdot,\cdot)$. When one statistical comparison is made, it is natural to reject the null hypothesis when pvalue is smaller than chosen $\alpha=0.05$. Since the goal of the experiment is to draw conclusions on the number of statistically different neighbors, we take into account Bonferroni correction. That is, we define the maximal pvalue that is still acceptable to reject the null hypothesis by $n^{2}=1024$. So, when the pvalue is smaller than $0.05/1024$, the corresponding points are considered different and the square is colored in blue in Figure 8. Otherwise, the null hypothesis is not rejected and the square is colored in red.
@@ 5.5 Distance-Driven UMDA
The global distribution in UMDA‑U does not tend to concentrate its probability mass on a few filters. As a result, when we sample a solution from this distribution, it often contains duplicates and exceeds the allowed number of distinct filters. To fix this, we introduce a second modification that adjusts sampling to respect the filter‐count constraint.
The idea of Probabilistic Logic Sampling, introduced for Bayesian Networks in [61], is applied in Estimation of Distribution algorithms that address permutations-based COPs [62]. We employ this idea to avoid repetitions of the same filter multiple times. Specifically, we create the conditional distribution $P(\mathrm{x}_{i}|\mathrm{x}_{1},\mathrm{x}_{2},\dots,\mathrm{x}_{i-1})$ to sample the component $\mathrm{x}_{i}$ of the candidate solution $\mathbf{x}$. If $\mathrm{x}_{i}\in\left\{\mathrm{x}_{1},\mathrm{x}_{2},\dots,\mathrm{x}_{i-1}\right\}$, then $P(\mathrm{x}_{i}|\mathrm{x}_{1},\mathrm{x}_{2},\dots,\mathrm{x}_{i-1})$ is set to 0, otherwise it is set to $C\cdot\mathbf{p}_{\mathrm{x}_{i}}$. Constant $C$ is chosen in such a way that: $\sum_{\mathrm{x}_{i}\in\mathbb{L}}P(\mathrm{x}_{i}|\mathrm{x}_{1},\mathrm{x}_{2},\dots,\mathrm{x}_{i-1})=1.$ We call this modification of UMDA-U as UMDA-U-PLS.
1For every $j\in\overline{1,L}$ initialize: $\mathrm{p}^{(1)}_{j}\leftarrow 1/L$;